OpenStudy (anonymous):

what is the solution to the equation sqrt 2x+8 -6=4

OpenStudy (anonymous):

\[\sqrt{2x+8} -6=4\]

OpenStudy (abb0t):

1. add 6 to both sides 2. square both sides 3. subtract 8 to both sides 4. divide by 2 the end.

OpenStudy (anonymous):

would it be x=\[\sqrt{5}\]

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

\[\sqrt{5}\]

OpenStudy (abb0t):

Now. How did you get 5?!

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

im confused

OpenStudy (abb0t):

Are you following the instructions I posted up above?

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

To square both sides means \((\sqrt{2x+8})\)\(\sf \color{red}{^2}\) = (10)\(\sf \color{red}{^2}\)

OpenStudy (anonymous):

my computer is not showing the formulas

OpenStudy (abb0t):

When you square both sides, the square root goes away.

OpenStudy (abb0t):

There should not even be a square root in your answer.

OpenStudy (anonymous):

would it be x=1

OpenStudy (abb0t):

No. How did you get 1?

OpenStudy (anonymous):

I squared both sides then subtracted by 8 and got 2x=2 then divided

OpenStudy (abb0t):

Go step by step. First, start by adding six to both sides: \(\sqrt{2x+8}\) + \(\sf \color{red}{6}\) - \(6\) = \(4 +6\) that gives you: \(\sf \color{blue}{\sqrt{2x+8}=10}\) NOW, square both sides.

OpenStudy (abb0t):

You square both sides so that you can solve for \(\sf \color{red}{x}\). You have it under the √ and in order to solve for it you MUST get rid of the √

OpenStudy (anonymous):

the formulas are not showing up on my computer so I cannot see what you typed

OpenStudy (anonymous):

would it be x=21?

OpenStudy (abb0t):

No.

OpenStudy (anonymous):

alright I give up I cant figure it out

OpenStudy (abb0t):

You're not following my directions. Square both sides. Remember, when you square something: \((10)\)\(\sf \color{red}{^2}\) = 100

OpenStudy (abb0t):

Also, \((\sqrt{a})^2 = a\)

OpenStudy (anonymous):

can you send it in that file again I cant see those things

OpenStudy (anonymous):

so x=46

OpenStudy (abb0t):

Yes! Correct.

OpenStudy (anonymous):

finally alright thank you

OpenStudy (abb0t):

Please practice.